DETERMINING REACTION MECHANISMS AND TRANSITION STATE GEOMETRIES IN LIQUIDS US-ING ONE AND TWO DIMENSIONAL ULTRAFAST VIBRATIONAL SPECTROSCOPY

\underline{JAMES F. CAHOON}, KARMA R. SAWYER, JACOB P. SCHLEGEL, MATTHEW C. ZOERB, AND CHARLES B. HARRIS, Department of Chemistry, University of California, Berkeley, CA 94720, and Chemical Sciences Division, Lawrence Berkeley National Laboratory, Berkeley, California 94720.

Both the thermal and photochemically driven reactions of organometallic complexes in solution can be readily examined with ultrafast one-dimensional (1D-IR) and two-dimensional infrared (2D-IR) vibrational spectroscopy. We show, for example, how the thermally-driven exchange of carbonyl groups in Fe(CO)_{5} (i.e. fluxionality) can be monitored with 2D-IR, providing direct evidence for the time-scale and transition-state of the exchange reaction. Applications of 1D and 2D-IR to photoinitiated chemical reactions will also be discussed.